What is the new molarity of the solution when water is added to 125.0 mL of 2.50 M NaOH to make 300. mL of a diluted NaOH solution?
1
Expert's answer
2020-11-22T11:11:44-0500

C1V1= C2V2

125×2.5=C2×300

C2 = 125×2.5/300 = 1.04M